OpenStudy (deoxna):

If x,y and z are the first three terms of an geometric sequence, show that x^2,y^2 and z^2 form another geometric sequence

OpenStudy (deoxna):

If anyone is interested, I solved it: U1= x U2= y = xr U3= z = xr^2 If the squares are geometric then (y^2/x^2)=(z^2/y^2) and of you work out the maths you get r^2=r^2, i.e. it is a sequence.